Jul 16, 2015 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11 a.m.

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Change BCAL and FCAL coordinates use the global coordinate system (not the target center). Action of this item is under discussion
  2. Understand BCAL efficiency estimates (cluster vs enhanced) for m7 and m8 data taking modes (Ahmed) Under active study
  3. Provide feedback to Will on shower fragmentation (All)
  4. Modify mcsmear to produce digihit outputs (Tegan) Completed and checked in.


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. BCAL Action Items and Working Tasks

Tentative Agenda

  1. Announcements
    • Time for meeting
  2. Action Items
  3. FCAL Update
  4. BCAL Update
  5. Calorimetry F250 Firmware Discussions
  6. Reconstruction and Simulation
  7. Calibration
  8. Any other business
  9. Testing


Attendees: David, Elton, Jon, Manuel, Mark, Adesh (JLab); Zisis, Ahmed (UofR); Sean (NW); Curtis, Will, Mike (CMU); George (Athens).

  1. Announcements
    • Start meetings at 11 am, including BCAL reconstruction meeting.
  2. Action Items
  3. FCAL Update
    1. Base / firmware (Jon)
      • Updated all firmware and testing on a single strand. Waiting for feedback from Hovanes, which will take about a week.
      • Will need to decide on a timeline and criteria for replacing poor performing bases. There are about 10 that need to be checked.
      • Before changing any hardware, the new firmware will be updated and problems verified.
    2. FCAL efficiency (Jon)
      • Starting to study exclusive omega p event with omega->pi0 gamma for determining FCAL efficiency
    3. Gain calibration. See Log Entry 3346084 (Adesh)
      • Gain calibrations (in ccdb) were determined by minimizing the pi0 width using B=0 runs
      • Nonlinearities (i.e. variation in pi0 mass with energy) were observed
      • Approximate nonlinearity corrections were added to ccdb, but were ineffective, so they have been backed out.
      • This calibration applied to B=not 0 runs result in a pi0 peak ~ 117 MeV, sigma~9 MeV.
      • An electron sample from the B=not 0 runs were used to determine gain constants for 20 < r < 60 cm, and used as a see for further pi0 width minimization. See Log Entry 3345736. The iteration used B=0 runs.
      • Output shows non-linearity for B=0 runs. However, non-linearity is not apparent in the B=not 0 runs.
      • Zisis: Suggest to limit range of fit of two-photon distribution to better reproduce the pi0 peak.
      • -> It appears that the B=0 and B=not 0 runs have different systematics
      • Eugene suggested yesterday to take LED data while ramping the magnet to assess the influence of B field on the pmt response.
      • Will: Suggests to start the pi0 width minimization with the pi0 peak centered on the physical mass. It could be that the fit is being because of the wrong overall normalization.
      • Elton: No pi0 peak is visible in the 500<E<600 MeV photon bin. Why not, and is it due to trigger thresholds. Need to understand this to better plan for more calibration runs next period.
  4. BCAL Update
    1. Time-walk (Zisis for Noemi)
      • Noemi has had some issues with generating timing plots. Versioning of gcc, Jana, etc have been resolved, but still nominal monitoring plots show a lack of entries. They are investigating if there is a problem with scale and events falling outside the nominal range.
      • Mark had sent early analysis of timing using LED data. Zisis asked if he used the unified hit quantities? Mark: Could not remember precisely, but thought he might have taken the data from DAQTree, but they should work with unified hit objects.
    2. BCAL efficiencies (Ahmed)
      • Working on understanding BCAL efficiencies. A modulation in phi has been attributed to inclusion of cosmic-rays into sample.
      • Trying to eliminate cosmic-ray events by cutting on energy in Fcal.
      • Mike: suggested some ways of eliminating cosmics and will send some code to do this. Elton: Using BCAL_Eff already uses matched tracks to showers, so most of the tracks should come from near the target, removing comic angular bias.
    3. Veff studies (George)
      • Work is ongoing, but George is planning to present an update at the next Cal Working meeting (in two weeks) with a summary and proposal for constants to be loaded into ccdb.
      • The current situation is that cosmics give a relatively uniform veff computation but require relatively large number of points to be excluded from fit and these are under investigation.
      • The cosmics also show a small but systematic difference between the top (~17 cm/ns) and bottom (~16.8 cm/ns), which is being studied.
  5. Calorimetry F250 Firmware Discussions
    • Summary (Elton)
    • David showed that small fraction of data is actually "content" (fadc, tdc info). Most of data are header, timing, etc words. Only 20% of BCAL data is content, and BCAl content is a few % of overall file size. At least for the moment, eliminating the pedestal word from the data will have a negligible effect on overall data volume
    • Mark did an LED study (with medium to large pulses) to understand the optimum integration window for the pulse "integral". The start of the window and the length of the window were varied. The conclusion was that for far side (no reflections), the optimum window was between 2 to 11 counts (relative to the threshold sample). Recall that currently we use -5 to 55 counts to cover the entire pulse. For near side pulses (includes reflection), the optimum is about 2 - 20 counts. Basically the beginning and end of the pulses contribute relatively more noise than the peak region. Studies with smaller pulses are still needed.
  6. Reconstruction and Simulation
    1. Early look at BCAL Clusters Log Entry 3346051.
      • Plotted "size" of clusters in Delta_t, Delta_z and Delta_phi.
      • Size is confined in time, but (on log plots), he finds that there are hits included into the clusters that have a wide range of angles phi and positions z.
      • Elton and Mark: Suggest to understand individual events that have wide distributions (e.g. with hdview2). Consider whether changes to cluster finder can address any problems.
      • Single ended hits contribute about 4% of energy at 1 GeV. Consider how and value of including these hits (in addition to double-ended points) to energy clusters.
    2. digiHits (Tegan)
      • Checked in digiHits, so that MC should now produced these objects.
      • Elton: The e-mail implied that with the new code the resulting MC showers would have a gain that was about 3 times too large.
      • Tegan agreed to update the variation=mc of gains to yield the (approximately) correct energy scale. [Note added: Already done. See e-mail]
  7. Calibration
  8. Any other business
  9. Testing